OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 3, – 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x − ___ Numerical Answers Expected! #INeedAnAnswer

OpenStudy (helder_edwin):

since x=-8 and x=2 then x+8=0 and x-2=0 ok?

OpenStudy (anonymous):

is that the anwser ?

OpenStudy (helder_edwin):

sorry \[ \large (x+8)(x-2)=x^2+6x-16 \]

OpenStudy (anonymous):

sorry wht ?

OpenStudy (helder_edwin):

now u have \[ \large y=x^2+6x-16 \] then \[ \large y+16+9=x^2+6x+9 \] \[ \large y+25=(x+3)^2 \] and this has vertex (-3,-25) as required

OpenStudy (anonymous):

sooo its -3,-25 ?

OpenStudy (helder_edwin):

16